Products are sold in batches of 20. A penalty clause requires that if any batch contains 2 or more defective items, the customer will receive a total refund on the cost of the batch.
How likely is this to happen if the proportion of defective products produced is 0.04?
Give your answer correct to 3 places of decimals.

Found 2 solutions by stanbon, Theo:
Answer by stanbon(75887)

Products are sold in batches of 20. A penalty clause requires that if any batch contains 2 or more defective items, the customer will receive a total refund on the cost of the batch.
How likely is this to happen if the proportion of defectives products produced is 0.04?
Give your answer correct to 3 places of decimals.
------------------
It's a binomial problem with n=20, p=0.04, 2<= x <= 20
P(2<= x <= 20) = 1 - binomcdf(20,0.04,1) = 0.1897
======================
Cheers,
Stan H.

Answer by Theo(13342)

probability of exactly 0 defective products in a batch is equal to .96^20 = .442002434

probability of exactly 1 defective product in a batch is equal to 20 * .04*.96^19 = .368335362

the probability that at least 2 will be defective is 1 minus the probability of (exactly 0 defective and exactly 1 defective).

This equals 1 - (.442002434 + .368335362) = .189662205.

If I did this correctly, the probability that at least 2 will be defective in a batch is 18.96%.